# definition

### v4

@@ 3 ID policy with reassignment
Once we obtain a solution to the LP in (5), we can construct a policy for each arm $i$, which we refer to as an optimal single-armed policy for arm $i$. In particular, let $(y^{*}_{i}(s,a))_{i\in[N],s\in\mathbb{S},a\in\mathbb{A}}$ be an optimal solution to the LP in (5). Then for arm $i$, the optimal single-armed policy, ${\bar{\pi}}^{*}_{i}$, is defined as
@@
The rationale behind considering these policies is as follows. If each arm $i$ individually follows its optimal single-armed policy ${\bar{\pi}}_{i}^{*}$, then the average reward per arm (total reward divided by $N$) achieves the upper bound $R_{N}^{\textup{rel}}$ given by the LP. However, this strategy only guarantees that the cost constraints are satisfied in a time-average sense, rather than conforming to the hard constraints in the original $N$-armed WCMDP. Thus, having each arm follow its optimal single-armed policy is not a valid policy for the original $N$-armed problem. Nevertheless, these optimal single-armed policies ${\bar{\pi}}_{i}^{*}$’s serve as guidance for how the arms should ideally behave to maximize rewards. The ID policy uses the ${\bar{\pi}}_{i}^{*}$’s as a reference. It is then designed to ensure that even under the hard cost constraints, most arms follow their optimal single-armed policies most of the time, yielding a diminishing gap to $R_{N}^{\textup{rel}}$ in reward.
@@
5:     $\text{newID}(i)\leftarrow i$ for all $i\in[N]$triangleright No need for ID reassignment
@@
7:     Initialize $\mathcal{F}\leftarrow\emptyset$triangleright Set of arms that have been assigned new IDs
@@
16:     For all $i\in[N]\setminus\mathcal{F}$, assign values to their $\text{newID}(i)$’s randomly from $[N]\setminus\{\text{newID}(i^{\prime})\colon i^{\prime}\in\mathcal{F}\}$
@@
We first define a few quantities that will be used in the ID reassignment algorithm. For each arm $i\in[N]$ and each cost type $k\in[K]$, the expected cost under the optimal single-armed policy is defined as $C_{k,i}^{*}=\sum_{s\in\mathbb{S},a\in\mathbb{A}}y_{i}^{*}(s,a)c_{k,i}(s,a)$. Based on $C_{k,i}^{*}$’s, we divide the cost constraints into active constraints and inactive constraints as follows. For each cost type $k\in[K]$, we say the type-$k$ cost constraint is active if
@@
and inactive otherwise. Let $\mathcal{A}\subseteq[K]$ denote the set of cost types corresponding to active constraints.
Based on the costs $C_{k,i}^{*}$’s and the active constraints, the ID reassignment algorithm rearranges arms so that the cost incurred by each contiguous segment of arms is approximately proportional to the length of the segment. We give a brief explanation of how the reassigned IDs affect the execution of the policy at the end of this section. The algorithm is formally described in Algorithm 1, with more details and properties provided in Appendix C. In the remainder of this paper, we use the reassigned IDs to refer to arms, i.e., arm $i$ refers to the arm whose new ID assigned by the ID reassignment algorithm is $i$.
@@
1:  Input: $N$-armed WCMDP instance $(\mathcal{M}_{i})_{i\in[N]}$
@@
We are now ready to describe our generalized ID policy, formally described in Algorithm 2. The policy begins with a one-time preprocessing phase: we solve the associated LP, construct the optimal single-armed policies, and reassign arm IDs using the ID reassignment algorithm (Algorithm 1). After the preprocessing, the policy proceeds at each time step $t$ as follows. For each arm $i$ (where $i$ is the reassigned ID), we first sample an action $\widehat{A}_{i,t}$, referred to as an ideal action, from the optimal single-armed policy ${\bar{\pi}}_{i}^{*}(\cdot\mid S_{i,t})$. We then attempt to execute these ideal actions, i.e., set the real actions equal to the ideal actions, in ascending order of arm IDs, starting from $i=1$, then $i=2$, and so on. We continue the attempt until we have used up at least one type of cost budget, at which point we let the remaining arms take action 00 (the no-cost action).
This ID policy is a natural generalization of the ID policy designed for homogeneous restless bandits [26]. At a high level, using arm IDs to decide the priority order for executing ideal actions guarantees that a subset of arms (those with smaller IDs) can persistently follow their ideal actions. This persistency gives these arms time to converge to their optimal state-action frequencies, which in turn enables the subset to gradually expand. In contrast, if we do not use IDs but instead randomly select a subset of arms to follow their ideal actions, the convergence may be disrupted. Indeed, [25] provides an example where this randomized strategy fails to achieve asymptotic optimality in restless bandits.
@@
In the heterogeneous setting, arms differ in their cost consumptions under their optimal single-armed policies. The ID reassignment algorithm is designed to prevent “plateaus” in cumulative cost as we progress from smaller to larger IDs. If such a plateau exists, the subset of arms allowed to follow their ideal actions (determined by the cost constraints) can become sensitive to the randomness in action sampling, potentially leading to performance instability. The ID reassignment algorithm ensures a regularity property of the cost consumptions, stated in Lemma 2 in Appendix C, which eliminates such plateaus.
4 Main result and technical overview
Before we present the main result, we first state our main assumption. This assumption is for the optimal single-armed policies ${\bar{\pi}}^{*}_{i}$’s. Note that each ${\bar{\pi}}^{*}_{i}$ is a stationary Markov policy. Therefore, under this policy, the state of arm $i$ forms a Markov chain. Let the transition probability matrix of this Markov chain be denoted as $P_{i}=(P_{i}(s,s^{\prime}))_{s\in\mathbb{S},s^{\prime}\in\mathbb{S}}$, where the row index is the current state $s$ and the column index is the next state $s^{\prime}$. Then $P_{i}(s,s^{\prime})$ can be written as
@@ 4 Main result and technical overview
Consider any $N$-armed WCMDP with initial system state $\bm{S}_{0}$ and assume that it satisfies Assumption 1. Let policy $\pi$ be the ID policy with reassignment (Algorithm 2). Then the optimality gap of $\pi$ is bounded as
@@
We re-emphasize that our proposed ID policy with reassignment is the first efficiently computable policy that achieves an $O(1/\sqrt{N})$ optimality gap for fully heterogeneous average-reward WCMDPs. In contrast, the best-known optimality gap for efficiently computable policies for average-reward WCMDPs is $o(1)$, achieved only under restrictive cost constraints and typed-heterogeneity.
@@
In the remainder of this section, we consider the ID policy, sometimes also referred to as policy $\pi$. Let ${\bm{X}}_{t}$ denote the system state under it, with the superscript $\pi$ omitted for brevity. To make this overview more intuitive, here let us assume that ${\bm{X}}_{t}$ converges to its steady state ${\bm{X}}_{\infty}$ in a proper sense such that taking expectations in steady state is the same as taking time averages. However, note that our formal results do not need this assumption and directly work with time averages. We call a function $V$ a Lyapunov/potential function if it maps each possible system state to a nonnegative real number.
@@
[…]
@@
Combining the above ideas, for any set of arms $D$, we let the Lyapunov function $h({\bm{X}}_{t},D)$ be the supremum of the differences between ${\bm{X}}_{t}$ and $\mu^{*}$ in all the features defined above with proper weightings:
@@
where $\gamma=\exp(-1/(2\tau))$ for $\tau$ defined in 1; each element $g\in\mathcal{G}$ is either $g=(r_{i}^{*})_{i\in[N]}$, or corresponds to the type-$k$ cost for some $k\in[K]$ (see Appendix D for the definition of $\mathcal{G}$). Note that dividing each term by powers of $\gamma$ is another interesting trick, which makes the Lyapunov function strictly contract under the optimal single-armed policies, as demonstrated in the proof of Lemma 3.
@@ Appendix A Detailed review on related work
Appendix B Proving the LP relaxation
In this section, we prove Lemma 1, which shows that the linear program in (5) is a relaxation of the WCMDP problem. Lemma 1 is restated as follows.
@@ Appendix B Proving the LP relaxation
Appendix C More details and properties of the ID reassignment algorithm
@@
The remaining budgets play a key role in the design of the ID reassignment algorithm. Roughly speaking, the goal of the ID reassignment is to ensure that when we expand a set of arms from $[n_{1}]$ to $[n_{2}]$ for some $n_{1}\leq n_{2}$, the drop in the remaining budget of any type $k$, i.e., $\overline{C}^{*}_{k}([n_{1}])-\overline{C}^{*}_{k}([n_{2}])$, is (almost) at least linear in $n_{2}-n_{1}$. Note that this property is automatically satisfied for $k$ if the type-$k$ constraint is inactive. This property is formalized in Lemma 2, and the need for it will become clearer in Appendix D when we introduce the so-called focus set in our analysis.
To achieve this desired property, we design our ID reassignment algorithm in the following way. If the set of active constraints is empty, i.e., $\mathcal{A}=\emptyset$, then there is no need to reassign the IDs. Otherwise, i.e., when $\mathcal{A}\neq\emptyset$, we first carefully choose two parameters, a positive real number $\delta$ and a positive integer $d$. We then divide the full ID set $[N]$ into groups of size $d$, i.e., $[d],[d+1:2d],[2d+1:3d],\dots,[(\lfloor N/d\rfloor-1)d+1:\lfloor N/d\rfloor d]$, and the remainder. We ensure that after the reassignment, each group contains at least one arm $i$ with $C_{k,i}^{*}\geq\delta$ for each active constraint type $k\in\mathcal{A}$.
@@
We are now ready to prove the inequality $\overline{C}^{*}_{k}([n_{1}])-\overline{C}^{*}_{k}([n_{2}])\geq\eta_{c}(n_{2}-n_{1})-M_{c}$ for any $n_{1},n_{2}$ with $1\leq n_{1}\leq n_{2}\leq N$. Consider the arms with new IDs in $[n_{1}:n_{2}]$. Let $g$ be the number of groups from groups of the form $\mathcal{I}(\ell)=[\ell d+1:(\ell+1)d]$ with $\ell=0,1,\dots,\lfloor N/d\rfloor-1$ that are completely contained in $[n_{1}:n_{2}]$. Then it is easy to see
@@
Appendix D Proof of main result (Theorem 1)

### v5

@@ 3 ID policy with reassignment
Once we obtain a solution to the LP in (5), we can construct a policy for each arm $i$, which we refer to as an optimal single-armed policy for arm $i$. In particular, let $(y^{*}_{i}(s,a))_{i\in[N],s\in\mathbb{S},a\in\mathbb{A}}$ be an arbitrary optimal solution to the LP in (5). Then for arm $i$, the optimal single-armed policy, ${\bar{\pi}}^{*}_{i}$, is defined as
@@
The rationale behind these policies is as follows. If each arm $i$ individually follows its optimal single-armed policy ${\bar{\pi}}_{i}^{*}$, then the average reward per arm (total reward divided by $N$) achieves the upper bound $R_{N}^{\textup{rel}}$ given by the LP. However, this strategy only guarantees that the budget constraints are satisfied in a time-average sense, rather than conforming to the hard constraints in the original $N$-armed WCMDP. Thus, having each arm follow its optimal single-armed policy is not a valid policy for the original $N$-armed problem. Nevertheless, these optimal single-armed policies ${\bar{\pi}}_{i}^{*}$’s serve as a guide for how the arms should ideally behave to maximize rewards. The ID policy uses the ${\bar{\pi}}_{i}^{*}$’s as a reference. It is then designed to ensure that even under the hard budget constraints, most arms follow their optimal single-armed policies most of the time, yielding a diminishing gap to $R_{N}^{\textup{rel}}$ in reward.
@@
5:  $\text{newID}(i)\leftarrow i$ for all $i\in[N]$ $\triangleright$ No need for ID reassignment
@@
7:  Initialize $\mathcal{F}\leftarrow\emptyset$ $\triangleright$ Set of arms that have been assigned new IDs
@@
16:  For all $i\in[N]\setminus\mathcal{F}$, assign values to their $\text{newID}(i)$’s randomly from $N]\setminus\{\text{newID}(i^{\prime})\colon i^{\prime}\in\mathcal{F}\}$
@@
We first define a few quantities that will be used in the ID reassignment algorithm. For each arm $i\in[N]$ and each cost type $k\in[K]$, the expected cost under the optimal single-armed policy is defined as $C_{k,i}^{*}=\sum_{s\in\mathbb{S},a\in\mathbb{A}}y_{i}^{*}(s,a)c_{k,i}(s,a)$. Based on $C_{k,i}^{*}$’s, we divide the budget constraints into active constraints and inactive constraints as follows. For each cost type $k\in[K]$, we say the type-$k$ budget constraint is active if
@@
and inactive otherwise. Let $\mathcal{A}\subseteq[K]$ denote the set of cost types corresponding to active constraints. Note that replacing $\alpha_{k}N/2$ with any constant and strict fraction of $\alpha_{k}N$ will not change the results.
Based on the costs $C_{k,i}^{*}$’s and the active constraints, the ID reassignment algorithm rearranges arms so that the cost incurred by each contiguous segment of arms is approximately proportional to the length of the segment. We give a brief explanation of how the reassigned IDs affect the execution of the policy at the end of this section. The algorithm is formally described in Algorithm 1, with more details and properties provided in Appendix D. In the remainder of this paper, we use the reassigned IDs to refer to arms, i.e., arm $i$ refers to the arm whose new ID assigned by the ID reassignment algorithm is $i$.
@@
1:Input: NN-armed WCMDP instance $(\mathcal{M}_{i})_{i\in[N]}$
@@
We are now ready to describe our generalized ID policy, formally described in Algorithm 2. The policy begins with a one-time preprocessing phase: we solve the associated LP, construct the optimal single-armed policies, and reassign arm IDs using the ID reassignment algorithm (Algorithm 1). After the preprocessing, the policy proceeds at each time step $t$ as follows. For each arm $i$ (where $i$ is the reassigned ID), we first sample an action $\widehat{A}_{i,t}$, referred to as an ideal action, from the optimal single-armed policy ${\bar{\pi}}_{i}^{*}(\cdot\mid S_{i,t})$. We then attempt to execute these ideal actions, i.e., set the real actions equal to the ideal actions, in ascending order of arm IDs, starting from $i=1$, then $i=2$, and so on. We continue the attempt until we have used up at least one type of cost budget, at which point we let the remaining arms take action 0 (the no-cost action).
This ID policy is a natural generalization of the ID policy designed for homogeneous restless bandits [26]. At a high level, using arm IDs to decide the priority order for executing ideal actions guarantees that a subset of arms (those with smaller IDs) can persistently follow their ideal actions. This persistency gives these arms time to converge to their optimal state-action frequencies, which in turn allows their instantaneous costs to converge to steady-state values. This convergence creates slack in the budget constraints, thereby allowing more arms to follow their ideal actions. In contrast, if we do not use IDs but instead randomly select a subset of arms to follow their ideal actions, the convergence may be disrupted. Indeed, [25] provides an example where this randomized strategy fails to achieve asymptotic optimality in restless bandits.
@@
In the heterogeneous setting, arms differ in their cost consumptions under their optimal single-armed policies. The ID reassignment algorithm is designed to prevent “plateaus” in cumulative cost as we progress from smaller to larger IDs. If such a plateau exists, the subset of arms allowed to follow their ideal actions (determined by the budget constraints) can become sensitive to the randomness in action sampling, potentially leading to performance instability. The ID reassignment algorithm ensures a regularity property of the cost consumptions, stated in Lemma 2 in Appendix D, which eliminates such plateaus.
4 Main results and technical overview
Before we present the main results, we first state our main assumption. This assumption is for the optimal single-armed policies ${\bar{\pi}}^{*}_{i}$’s. Note that each ${\bar{\pi}}^{*}_{i}$ is a stationary Markov policy. Therefore, under this policy, the state of arm $i$ forms a Markov chain. Let the transition probability matrix of this Markov chain be denoted as $P_{i}=(P_{i}(s,s^{\prime}))_{s\in\mathbb{S},s^{\prime}\in\mathbb{S}}$, where the row index is the current state $s$ and the column index is the next state $s^{\prime}$. Then $P_{i}(s,s^{\prime})$ can be written as
@@ 4 Main results and technical overview
Next, we state our main theorem, Theorem˜1, whose proof is provided in Appendix˜E.
@@
Consider an NN-armed WCMDP problem satisfying Assumption 1, with initial system state $\bm{S}_{0}$. Let policy $\pi$ be the ID policy with reassignment (Algorithm 2). Then the optimality gap of $\pi$ is bounded as
@@
We re-emphasize that our proposed ID policy with reassignment is the first efficiently computable policy that achieves an $O(1/\sqrt{N})$ optimality gap for fully heterogeneous average-reward WCMDPs. In contrast, the best-known optimality gap for efficiently computable policies for average-reward WCMDPs is $o(1)$, achieved only under restrictive budget constraints and typed-heterogeneity.
@@
Remark 1 (Generalization of result).
Our result can be generalized to the setting where a $g(N)$ fraction of arms have unbounded mixing times, and the mixing times of the remaining arms scale with $N$ with an upper bound $\tau$. In this case, we can modify the ID policy by reassigning this $g(N)$ fraction of arms the largest IDs, effectively ignoring these arms. Applying Theorem˜1 to the remaining arms then implies $R^{*}(N,\bm{S}_{0})-R(\pi,\bm{S}_{0})\leq O\big(K^{5}\max\{r_{\max},c_{\max}\}^{7}\tau^{4}/(\alpha_{\min}^{6}\sqrt{N})+r_{\max}g(N)\big)$. Consequently, this modified policy is asymptotically optimal when $g(N)=o(1)$ and $\tau=o(N^{1/8})$.
Adapting the proof of Theorem˜1 also gives the following finite-time bound (see Appendix˜I).
Proposition 1 (Finite-time bound).
Under the same conditions as Theorem˜1, we have for any $T\geq 1$,
$R^{*}(N,\bm{S}_{0})-\frac{1}{TN}\sum_{t=0}^{T-1}\sum_{i\in[N]}\mathbb{E}\left[r_{i}(S_{i,t}^{\pi},A_{i,t}^{\pi})\right]\leq\frac{C_{\textnormal{ID}}}{\sqrt{N}}+\frac{C_{\textnormal{finite}}}{T},$ (12)
where $C_{\textnormal{finite}}$ is another positive constant independent of $N$.
@@
In the remainder of this section, we consider the ID policy, also referred to as the policy $\pi$. Let ${\bm{X}}_{t}$ denote the system state under it, with the superscript $\pi$ omitted for brevity. To make this overview more intuitive, here let us assume that ${\bm{X}}_{t}$ converges to its steady state $\bm{X}$$_{\infty}$ in a proper sense such that taking expectations in steady state is the same as taking time averages. However, note that our formal results do not need this assumption and directly work with time averages. We call a function $V$ a Lyapunov/potential function if it maps each possible system state to a nonnegative real number.
@@
where ${\bm{X}}_{t}(D)=(X_{t}(D,s))_{s\in\mathbb{S}}$ denotes within $D$, the number of arms in each state $s$, divided by $N$; $m(D)=|D|/N$; and the norm $\left\lVert\cdot\right\rVert$ is a properly defined norm. The idea is that if all arms in $D$ follow the optimal single-armed policy, the state distribution of each arm in $D$ gets closer to $\mu^{*}$, and thus ${\bm{X}}_{t}(D)$ gets closer to $m(D)\mu^{*}$ over time.
In the heterogeneous setting, we also want to construct a Lyapunov function $h({\bm{X}}_{t},D)$ to witness the convergence of any set of arms $D$ if they follow the optimal single-armed policies. However, unlike the homogeneous setting, now it no longer makes sense to aggregate arm states into an empirical state distribution, since each arm’s dynamics is distinct. Instead, our Lyapunov function considers $X_{i,t}-\mu_{i}^{*}$, where recall $X_{i,t}(s)$ is the indicator that arm $i$’s state is $s$ at time $t$. A naive first attempt is to construct the Lyapunov function from the pointwise distances, $\left\lVert X_{i,t}-\mu_{i}^{*}\right\rVert$ for each arm $i$, with a properly defined norm $\left\lVert\cdot\right\rVert$. However, the pointwise distances are very noisy: $\left\lVert X_{i,t}-\mu_{i}^{*}\right\rVert$ could be large even when the state of arm $i$ independently follows the distribution $\mu_{i}^{*}$ for each $i$, a situation when we should view the set of arms as already converged.
@@
Combining the above ideas, for any set of arms $D$, we let the Lyapunov function $h({\bm{X}}_{t},D)$ be the supremum of the differences between ${\bm{X}}_{t}$ and $\mu^{*}$ in all the features directions defined above, under proper weightings:
@@
where $\gamma=\exp(-1/(2\tau))$ for $\tau$ defined in ˜1; each element $g\in\mathcal{G}$ is either $g=(r_{i}^{*})_{i\in[N]}$, or corresponds to the type-$k$ cost for some $k\in[K]$ (See Appendix˜E for the definition of $\mathcal{G}$). Note that dividing each term by powers of $\gamma$ is another interesting trick, which induces a negative drift in $h({\bm{X}}_{t},D)$ under the optimal single-armed policies, as demonstrated in the proof of Lemma˜3.
@@
5 Experiments
Figure 1: Asymptotic optimality of ID policy.
Figure 2: ID policy vs ERC policy.
In this section, we perform two sets of experiments to illustrate the numerical performance of the proposed ID policy for fully heterogeneous WCMDPs.
In the first set of experiments, we demonstrate the asymptotic optimality of the ID policy. We increase the number of arms as $N\in\{100,200,400,800,1600,3200\}$. Each arm’s MDP has $10$ states and $4$ actions, with parameters generated uniformly at random in a proper sense. The $N$-armed problem has $4$ budget constraints, with cost functions also generated randomly. More details are provided in Section˜B.1. We simulate the policy for $2\times 10^{4}$ time steps over $4$ replications for each $N$. To illustrate the performance more clearly, we measure the optimality ratio, defined as the ratio between the long-run average reward achieved by a policy and the LP relaxation upper bound $R^{\text{rel}}_{N}$. Confidence intervals are calculated using the batch means method with a batch size of $4000$, but they are typically too small to be visible on figures. Figure 2 shows that the optimality ratio of the ID policy becomes increasingly close to $1$ as $N$ increases.
In the second set of experiments, we compare the ID policy with the ERC policy proposed in [46]. As discussed in Section˜1, only a few prior papers address heterogeneous WCMDPs. Among them, [46] considers the most general setting, but is still limited to a single-budget constraint, state-independent costs, and typed heterogeneity. To make a fair comparison, we evaluate both policies under this special case, while keeping all other settings the same as in the first set of experiments. ID policy turns out to have a slight improvement over ERC policy in some instances, one of which is shown in Figure 2, and has comparable performance in others. Importantly, unlike ERC, the ID policy applies to far more general classes of WCMDPs.
@@ Appendix A Detailed review on related work
Appendix B Experimental details
In this appendix, we provide details of the two WCMDP instances considered in Section˜5, and the definition of the baseline policy, the ERC policy from [46]. The complete code for these experiments is available on GitHub [53], and all results can be reproduced within 24 hours on a standard PC (e.g., 6-Core Intel Core i7).
B.1 WCMDP instance generation
Details of the WCMDP instance 1 (In Figure 2).
$|\mathbb{S}|=10$, $|\mathbb{A}|=4$, $K=4$. For each $i\in[N]$, $s\in\mathbb{S}$, and $a\in\mathbb{A}$, $r_{i}(s,0)=0$, and $r_{i}(s,a)$ is independently sampled from the uniform distribution over $[0,1]$ for each $a\neq 0$; $P_{i}(s,a,\cdot)$ is independently sampled from the uniform distribution over the probability simplex. As for the cost function, for each $i\in[N]$, $s\in\mathbb{S}$, $a\in\mathbb{A}$, and $k\in[K]$, we have $c_{k,i}(s,0)=0$, and $c_{k,i}(s,a)$ is independently sampled from the uniform distribution over $[0,1]$ for each $a\neq 0$. For each $k\in[K]$, $\alpha_{k}$ is uniformly sampled from $\{0.05,0.1,0.15,\dots,0.45\}$, i.e., the uniform distribution over integer multiples of $0.05$ in the interval $(0,0.5)$.
Details of the WCMDP instance 2 (In Figure 2).
This instance is typed heterogeneous with $10$ types, with equal fraction of arms in each type. Each arm has $|\mathbb{S}|=10$ states, $|\mathbb{A}|=4$ actions. For each type, the reward function and transition kernel are generated in the same way as each arm in ithe first instance. The budget is also sampled from the same distribution as the first instance. There is a single budget constraint. The cost function depends only on the action, consistent with the setting in [46]; the cost of each of action $a\in\mathbb{A}$ is sampled from the uniform distribution over $[0,1]$.
B.2 Definition of ERC policy
The ERC policy [46] solves the same LP in (5) to define the single-armed policies ${{\bar{\pi}}^{*}}_{i}(a|s)$ as in (7). At each time step, the policy computes an index for each arm-state pair $(i,s)\in[N]\times\mathbb{S}$:
$\mathcal{I}(i,s)=\sum_{a\in\mathbb{A}}{{\bar{\pi}}^{*}}_{i}(a|s)r_{i}(s,a).$
It then iterates through the arms in descending order of these indices. For each arm, it samples an action $a\sim{{\bar{\pi}}^{*}}_{i}(\cdot\mid s)$ and applies it only if the budget is sufficient. If not, it defaults to action 0.
Appendix C Proving the LP relaxation
In this section, we prove Lemma 1, which shows that the linear program in (5) is a relaxation of the WCMDP problem. Lemma˜1 is restated as follows.
@@ Appendix C Proving the LP relaxation
Appendix D More details and properties of the ID reassignment algorithm
@@
The remaining budgets play a key role in the design of the ID reassignment algorithm. Roughly speaking, the goal of the ID reassignment is to ensure that when we expand a set of arms from $n_{1}]$ to $n_{2}]$ for some $n_{1}\leq n_{2}$, the drop in the remaining budget of any type $k$, i.e., $\overline{C}^{*}_{k}([n_{1}])-\overline{C}^{*}_{k}([n_{2}])$, is (almost) at least linear in $n_{2}-n_{1}$. Note that this property is automatically satisfied for $k$ if the type-$k$ constraint is inactive. This property is formalized in Lemma 2, and the need for it will become clearer in Appendix˜E when we introduce the so-called focus set in our analysis.
To achieve this desired property, we design our ID reassignment algorithm in the following way. If the set of active constraints is empty, i.e., $\mathcal{A}=\emptyset$, then there is no need to reassign the IDs. Otherwise, i.e., when $\mathcal{A}\neq\emptyset$, we first carefully choose two parameters, a positive real number $\delta$ and a positive integer $d$. We then divide the full ID set $[N]$ into groups of size $d$, i.e., $d],[d+1:2d],[2d+1:3d],\dots,[(\lfloor N/d\rfloor-1)d+1:\lfloor N/d\rfloor d]$, and the remainder. We ensure that after the reassignment, each group contains at least one arm $i$ with $C_{k,i}^{*}\geq\delta$ for each active constraint type $k\in\mathcal{A}$.
@@
We are now ready to prove the inequality $\overline{C}^{*}_{k}([n_{1}])-\overline{C}^{*}_{k}([n_{2}])\geq\eta_{c}(n_{2}-n_{1})-M_{c}$ for any $n_{1},n_{2}$ with $1\leq n_{1}\leq n_{2}\leq N$. Consider the arms with new IDs in $n_{1}:n_{2}]$. Let $g$ be the number of groups from groups of the form $\mathcal{I}(\ell)=[\ell d+1:(\ell+1)d]$ with $\ell=0,1,\dots,\lfloor N/d\rfloor-1$ that are completely contained in $n_{1}:n_{2}]$. Then it is easy to see
@@
Appendix E Proof of main result (Theorem 1)